### v1

@@ 3 Proofs of the main theorems
The Postnikov completion $(-)^{\leq\bullet}:\mathcal{E}\to\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$ and the hypercompletion $h^{*}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$ invert the same class of maps, namely the $\infty$/̄connected maps. If all Postnikov towers are effective and so $(-)^{\leq\bullet}:\mathcal{E}\to\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$ is a localization, then the canonical morphism $\mathcal{E}^{\leq\infty}\to\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$ is an equivalence. When $\mathcal{E}$ satisfies $\mathsf{CC}_{d}$ for some finite $d$, this holds by Lemmas 3.1 and 3.4. ∎
@@
Recall from 2.20 (e) that every hypercomplete topos satisfies $\mathsf{CC}_{\infty}$. Since there exist hypercomplete topoi which are not Postnikov complete, this shows that Theorem 1.1 is false for $d=\infty$.
@@
Recall from [Lur18, Corollary A.7.2.6] that the subcategory of Postnikov complete $\infty$/̄topoi is closed under colimits of geometric morphisms. The $\infty$/̄topos of sheaves on the cube $[0,1]^{d}$ is Postnikov complete because it is locally of homotopy dimension $\leq d$ [Lur09, Proposition 7.2.1.10]. The $\infty$/̄topos $\coprod_{d}\,[0,1]^{d}$ of 2.20 (g) is Postnikov complete as a colimit of Postnikov complete $\infty$/̄topoi. This shows that the converse of Theorem 1.1 is false. Another counterexample is given by the $\infty$/̄topos of Proposition 4.2.
@@
By putting together Theorems 1.1 and 1.2 with 2.29 (c), we get a proof of [MR22, Theorem A].
@@ Appendix A Postnikov separation, convergence, effectivity, and completeness
where the first functor is a localization and the second functor is conservative. [^5] The factorization is typical of separation–completion processes: for example, the separation–completion of a ring $A$ at an ideal $I$: $A\to A/\bigcap I^{n}\to\lim A/I^{n}$, or the separation–completion of a pseudometric space $(X,d)$: $X\to X/(d=0)\to\overline{X}$. For this reason, we suggest replacing the terminology hypercompletion by Postnikov separation. The truncations of an object $X$ and their limit fit into a diagram
@@
Remark A.4.
A similar terminology is used for prestable categories in [Lur18, Definition C.1.2.12].
@@
Most of the statements are reformulations of Definitions A.2 and A.3. We only give an arguments for (4). If formal Postnikov towers are effective, the functor $\mathcal{C}\to\mathrm{\mathcal{P}ost}\left(\mathcal{C}\right)$ is a localization and is therefore an equivalence if and only if it is fully faithful or conservative. ∎

### v2

@@ 3 Proofs of the main theorems
The Postnikov completion $(-)^{\leq\bullet}:\mathcal{E}\to\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$ and the hypercompletion $P_{\infty}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$ invert the same class of maps, namely the $\infty$/̄connected maps. If all Postnikov towers are effective and so $(-)^{\leq\bullet}:\mathcal{E}\to\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$ is a localization, then the canonical morphism $\mathcal{E}^{\leq\infty}\to\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$ is an equivalence. When $\mathcal{E}$ satisfies $\mathsf{CC}_{d}$ for some finite $d$, this holds by Lemmas 3.1 and 3.4. ∎
@@
By Theorems 1.1 and 2.27, if $\mathsf{CC}_{d}$ holds in $\mathcal{E}$, it holds also in $\mathrm{\mathcal{P}ost}\left(\mathcal{E}\right)$.
@@
Recall from 2.20 (f) that every hypercomplete topos satisfies $\mathsf{CC}_{\infty}$. Since there exist hypercomplete topoi which are not Postnikov complete, this shows that Theorem 1.1 is false for $d=\infty$.
Remark 3.8.
Recall from [Lur18, Corollary A.7.2.6] that the subcategory of Postnikov complete $\infty$/̄topoi is closed under colimits of geometric morphisms. The $\infty$/̄topos of sheaves on the cube $[0,1]^{d}$ is Postnikov complete because it is locally of homotopy dimension $\leq d$ [Lur09, Proposition 7.2.1.10]. The $\infty$/̄topos $\coprod_{d}\,[0,1]^{d}$ of 2.20 (h) is Postnikov complete as a colimit of Postnikov complete $\infty$/̄topoi. This shows that the converse of Theorem 1.1 is false. Another counterexample is given by the $\infty$/̄topos of Proposition 4.2.
Remark 3.9.
By 2.20 (b), the conclusion of Theorem 1.1 holds for all slices of $\mathcal{E}$. We do not know if the slices of Postnikov complete topoi are always Postnikov complete.
@@
By putting together Theorems 1.1 and 1.2 with 2.35 (c), we get a proof of [MR22, Theorem A].
@@ Appendix A Postnikov separation, convergence, effectivity, and completeness
where the first functor is a localization and the second functor is conservative. The truncations of an object $X$ and their limit fit into a diagram
@@
Remark A.1.
The factorization $\mathcal{C}\to\mathcal{C}^{\leq\infty}\to\mathrm{\mathcal{P}ost}\left(\mathcal{C}\right)$ is typical of separation–completion processes. Other examples includes the separation–completion $A\to A/\bigcap I^{n}\to\lim A/I^{n}$ of a ring $A$ at an ideal $I$, or the separation–completion $X\to X/(d=0)\to\overline{X}$ of a pseudometric space $(X,d)$. For this reason, we suggest replacing the terminology hypercompletion by Postnikov separation. A similar terminology is used for prestable categories in [Lur18, Definition C.1.2.12].
@@
Most of the statements are reformulations of Definitions A.3 and A.4. We only give an argument for (4). If formal Postnikov towers are effective, the functor $\mathcal{C}\to\mathrm{\mathcal{P}ost}\left(\mathcal{C}\right)$ is a localization and is therefore an equivalence if and only if it is fully faithful or conservative. ∎